# Attitude Trajectories Stabilizable by Linear Control

### v1

@@ Abstract
It is quite often claimed, and correctly so, that linear methods cannot achieve global stability results for attitude control, and conversely that nonlinear control is essential in order to achieve (almost) globally stable tracking of general attitude trajectories. On account of this definitive result, and also because of the existence of powerful nonlinear control techniques, there has been relatively very little work analyzing the limits and performance of linear attitude control. It is the purpose of this paper to provide a characterization of the stability achievable for one class of linear attitude control problems, namely those leading to a constant quaternion difference. In this paper, we analytically derive a critical error angle below which linearized dynamics lead to natural marginal stability for such a system, and above which the system is unstable. The dynamics are then used to derive a locally stable linear attitude controller whose performance is validated using simulations.
@@ I Introduction
The attitude of a vehicle may be controlled to track a desired trajectory beginning from an (almost) arbitrary initial condition using a variety of (almost) globally stable nonlinear control algorithms, such as [1][2]. The controller in [1] uses a quaternion error variable to derive the controller, and this is extended in several works such as [3] to axisymmetric bodies, in [4] for neural network applications, and in [5] to account for modelled dissipation. An orthogonal matrix based representation of the attitude error is used in [6] to achieve similar stability results, and has been extended in several works such as [7] for robust control, [8] for output-feedback attitude control, [9] for flexibile bodies, and [2] for constrained systems.
Despite the known limitations of linear attitude control, its use is still widespread on account of the associated simplicity, and linear response to principal angle errors. For example, [10] used a linearization of the double integral form of attitude dynamics to regulate large initial errors. A controller that is linear in terms of the quaternion vector component is described in [11], which has the associated limitation of windup when the quaternion error angle is greater than 180 degrees. Linear PD and LQR methods for controlling a spacecraft’s Euler-angle attitude are compared in [12] leading to the conclusion that LQR methods lower the required control effort for similar attitude errors. In [13], the authors describe a linear attitude controller using Euler angle errors for a quadrotor UAV, and the controller works as long as the quadrotor’s attitude is sufficiently far away from the gimbal-lock singularity. In fact, such linear controllers are still prevalent in many commercial jet airplane autopilots which are for the most part far away from the singular Euler angle attitude. Other similar examples from literature include a design using the truncated Taylor series approximation for the quaternion exponentials in order to achieve a fast single-axis reorientation slew-rate on agile spacecrafts in [14], a PID attitude controller in [15], and an optimal linear controller in [16].
In this paper, we are interested in analyzing the stability of linear quaternion differential equations in time. In particular we consider the special case of a constant quaternion difference, since the results for this special case are sufficiently novel (the results have not been published previously in the literature to the best of the authors’ knowledge) and remarkable (being one of the few problems whose solutions can be analyzed and solved by hand). The chief contributions of this article may be summarized by the conditions provided on the error angle and the quaternion vector error in (14) and (15) for ensuring constant difference constant angle attitude tracking, the characteristic polynomial (28) for the linearized system, and the discriminant in (30) which determines marginal stability.
@@ II Notation and Terminology
The attitude of a rigid body will be represented by a unit-magnitude 4-component quaternion $\check{q}=[q_{0}\quad q_{1}\quad q_{2}\quad q_{3}]^{T}$, with a check accent above the symbol representing it. The first component $q_{0}$ is also referred to as the scalar component, while the three components $[q_{1}\quad q_{2}\quad q_{3}]^{T}$ are also referred to as the vector part $\vec{q}_{v}$ of $\check{q}$. The most important law of quaternion algebra is their multiplication rule [17]:
$\check{p}\otimes\check{q}$ $=\begin{bmatrix}p_{0}\\ \vec{p}_{v}\end{bmatrix}\otimes\begin{bmatrix}q_{0}\\ \vec{q}_{v}\end{bmatrix}=\begin{bmatrix}p_{0}q_{0}-\vec{p}_{v}^{T}\vec{q}_{v}\\ p_{0}\vec{q}_{v}+q_{0}\vec{p}_{v}+\vec{p}_{v}\times\vec{q}_{v}\end{bmatrix},$ (1)
where $\vec{p}_{v}\times\vec{q}_{v}$ (third term in the vector part of the product) denotes the usual vector product of two Euclidean vectors in three-dimensional Euclidean space. The skew-symmetric matrix corresponding to the vector cross product will be denoted as $[\vec{p}_{v}\times]$ so that $\vec{p}_{v}\times\vec{q}_{v}$ may also be expressed in matrix notation as $[\vec{p}_{v}\times]\vec{q}_{v}$. It follows from the above multiplication rule that the multiplicative identity is the quaternion with unit scalar part and zero vector part, and that the inverse of a unit-magnitude quaternion is given by negating its vector part [17]:
$\check{q}$ $=\begin{bmatrix}q_{0}\\ \vec{q}_{v}\end{bmatrix}\Rightarrow\check{q}^{-1}=\begin{bmatrix}q_{0}\\ -\vec{q}_{v}\end{bmatrix}.$ (2)
@@
It may be recalled that unit quaternions provide a minimal, singularity-free, and global description of rotations [18], with composition of rotations achieved by multiplication of quaternions. The kinematics of an attitude quaternion may be expressed in terms of the body-frame angular velocity $\vec{\omega}$ as [17]
$\dot{\check{q}}$ $=\frac{1}{2}\check{q}\otimes\vec{\omega}.$ (3)
When Euclidean 3-vectors such as $\vec{\omega}$ are used as quaternion operands as in (3), they are prefixed with a scalar component of zero.
@@
$\vec{n}$ $=\vec{\omega}\times J\vec{\omega}+J\dot{\vec{\omega}},$ (4)
where $\vec{n}\in\mathbb{R}^{3}$ is the net external torque, and $J\in\mathbb{R}^{3\times 3}$ is the (second) moment of inertia of the rigid body in the body-fixed coordinate frame. Assuming a solid body, the moment of inertia is positive definite, and the above equation may be inverted to obtain a one-to-one correspondence between the angular acceleration and the applied external torque. Therefore, one may justifiably consider the angular acceleration itself as the user input instead of the torque.
@@
$\dot{\check{p}}$ $=\frac{1}{2}\check{p}\otimes\vec{\nu},$ (5)
where $\vec{\nu}$ is the reference angular velocity. We will assume that the angular acceleration $\dot{\vec{\nu}}$ is also given to us. We will sometimes use the half-angular-velocities, $\vec{v}$ for $\vec{\nu}/2$ and $\vec{w}$ for $\vec{\omega}/2$, to avoid the factors of 2 and reduce notational clutter. In terms of $\vec{v}$ and $\vec{w}$, we have
$\dot{\check{p}}$ $=\check{p}\otimes\vec{v},\;\dot{\check{q}}=\check{q}\otimes\vec{w}.$ (6)
III Linear constant difference attitude control
@@
denote the constant difference between the desired attitude $\check{p}$ and the actual attitude $\check{q}$. The difference quaternion $\check{r}$ is clearly not a unit quaternion in general, but that does not preclude linear control based upon it.
We are therefore interested in studying the linear (small perturbation) stability of solutions to the quaternion differential equation
$\dot{\check{q}}=\dot{\check{p}}.$ (8)
Define the error quaternion as
@@
with (6) yielding the dynamics
$\dot{\check{e}}$ $=-\vec{v}\otimes\check{p}^{-1}\otimes\check{q}+\check{p}^{-1}\otimes\check{q}\otimes\vec{w}$
$=\check{e}\otimes\vec{w}-\vec{v}\otimes\check{e}.$ (10)
The reader may recall our convention that pure vectors are prefixed with a scalar component of zero when used as quaternion operands. So $\vec{w}$ in $\check{e}\otimes\vec{w}$ is really the 4-vector $[0\quad\vec{w}^{T}]^{T}$.
@@
$e_{0}$ $=p_{0}q_{0}+\vec{p}_{v}^{T}\vec{q}_{v}=p_{0}(p_{0}+r_{0})+\vec{p}_{v}^{T}(\vec{p}_{v}+\vec{r}_{v})$
$=1+r_{0}p_{0}+\vec{r}_{v}^{T}\vec{p}_{v}.$ (11)
@@
Since the right hand side of the above equation is a constant, we see that $e_{0}$, and therefore the error angle, are also constant in time.
Using (6) in (8), we obtain the mathematical statement which encapsulates the constant difference constraint as a relationship between the desired and actual (half) angular velocities, $\vec{v}$ and $\vec{w}$:
$\vec{v}$ $=\check{e}\otimes\vec{w},$
@@
$\vec{e}_{v}^{T}\vec{w}$ $=0,$ (14)
$\vec{v}$ $=e_{0}\vec{w}+\vec{e}_{v}\times\vec{w}.$ (15)
@@
It may be noted that we could have as well used the constant difference constraint and the inverse of (9), to infer that $\vec{v}$ is also perpendicular to $\vec{e}_{v}$ like $\vec{w}$, and to express $\vec{w}$ in terms of $\vec{v}$ as follows
$\vec{w}=\check{e}^{-1}\otimes\vec{v}\;\Rightarrow\vec{e}_{v}^{T}\vec{v}=0,\;\vec{w}=e_{0}\vec{v}-\vec{e}_{v}\times\vec{v}.$ (16)
The scalar equation (14) implies that the body (half) angular velocity $\vec{w}$ must be perpendicular to the axis of the error rotation $\vec{e}_{v}$ in order to sustain a constant difference between $\check{p}$ and $\check{q}$. The relationship in (15) may be utilized to eliminate the desired (half) angular velocity $\vec{v}$ in (15) and express the kinematics of the error quaternion solely in terms of its own components $e_{0}$ and $\vec{e}_{v}$ and the actual (half) angular velocity $\vec{w}$:
$\dot{\check{e}}$ $=\begin{bmatrix}\vec{e}_{v}^{T}(\vec{v}-\vec{w})\\ e_{0}(\vec{w}-\vec{v})+\vec{e}_{v}\times(\vec{v}+\vec{w})\end{bmatrix}$
$=\begin{bmatrix}(e_{0}-1)\vec{e}_{v}^{T}\vec{w}\\ (e_{0}-1)\vec{w}+\vec{e}_{v}\times\vec{w}+\vec{e}_{v}\vec{e}_{v}^{T}\vec{w}\end{bmatrix},$ (17)
where we have expanded $\vec{v}$ in terms of $\vec{w}$ using (15).
We have so far utilized the relation in (15) to derive the kinematic governing equation (17). The scalar part of (17) simply yields the result derived earlier in (13), that the scalar component $e_{0}$ of the error quaternion, or equivalently the error angle, remains a constant when the angular velocity is perpendicular to the axis of the error rotation.
@@
The vector part of (17) effectively contains three scalar equations in terms of six scalar components, three each in the vectors $\vec{e}_{v}$ and $\vec{w}$. In order to close the system, we need to bring in attitude kinetics and derive an equation for the body (half) angular acceleration $\dot{\vec{w}}$ which invokes the control torques and moments through (4).
While the kinematic equation (17) for $\dot{\vec{e}}_{v}$ cannot be simplified any further without additional assumptions, the kinetic equation allows us some liberty at attempting to achieve the constraint in (14). Taking the derivative of the scalar equation (14), and using (17), we obtain one scalar equation:
$\vec{e}_{v}^{T}\dot{\vec{w}}=-\vec{w}^{T}\dot{\vec{e}}_{v}=(1-e_{0})|\vec{w}|^{2}-(\vec{e}_{v}^{T}\vec{w})^{2}.$ (18)
It is worth emphasizing that although (18) does not explicitly invoke the desired angular acceleration $\dot{\vec{v}}$, the equation does in fact impose one scalar degree of constraint on $\dot{\vec{v}}$ on account of the relationship in (15). We may now design a minimal $\dot{\vec{w}}$ with no components along any other direction but $\vec{e}_{v}$:
$\dot{\vec{w}}$ $=\frac{(1-e_{0})|\vec{w}|^{2}\vec{e}_{v}}{|\vec{e}_{v}|^{2}}=\frac{|\vec{w}|^{2}\vec{e}_{v}}{1+e_{0}}.$ (19)
The two remaining degrees of freedom in $\dot{\vec{w}}$, namely the other two components along $\vec{w}$ and $\vec{e}_{v}\times\vec{w}$, may be passed on to designing $\dot{\vec{v}}$ without compromising what follows in this section. (See Section V for one possible linear design that yields asymptotic convergence on the angular velocity.)

### v2

# Linear Stability Analysis of a Constant Quaternion Difference Attitude Controller

@@ Abstract
It is quite often claimed, and correctly so, that linear methods cannot achieve global stability results for attitude control, and conversely that nonlinear control is essential in order to achieve (almost) globally stable tracking of general attitude trajectories. On account of this definitive result, and also because of the existence of powerful nonlinear control techniques, there has been relatively very little work analyzing the limits and performance of linear attitude control. This paper provides a characterization of the stability achievable for one class of linear attitude control problems, namely those leading to a constant quaternion difference, and shows how the strategy can be used to design an efficient attitude controller. In this paper, we analytically derive a critical error angle below which linearized dynamics lead to natural marginal stability for such a system, and above which the system is unstable. The dynamics are then used to derive a locally stable linear attitude controller whose performance is validated using simulations.
@@ I Introduction
The attitude of a vehicle may be controlled to track a desired trajectory beginning from an (almost) arbitrary initial condition using a variety of (almost) globally stable nonlinear control algorithms, such as [wen:91a][shi:23a]. The controller in [wen:91a] uses a quaternion error variable to derive the controller, and this is extended in several works such as [tsiotras:95a] to axisymmetric bodies, in [fortuna:01a] for neural network applications, and in [mitikiri:22a] to account for modelled dissipation. An orthogonal matrix based representation of the attitude error is used in [chaturvedi:11a] to achieve similar stability results, and has been extended in several works such as [lee:13a] for robust control, [zlotnik:14a] for output-feedback attitude control, [chen:19a] for flexibile bodies, and [shi:23a] for constrained systems.
Despite the known limitations of linear attitude control, its use is still widespread on account of the associated simplicity, and linear response to principal angle errors. For example, [dwyer:84a] used a linearization of the double integral form of attitude dynamics to regulate large initial errors. A controller that is linear in terms of the quaternion vector component is described in [bach:93a], which has the associated limitation of windup when the quaternion error angle is greater than 180 degrees. Linear PD and LQR methods for controlling a spacecraft’s Euler-angle attitude are compared in [beatty:06a] leading to the conclusion that LQR methods lower the required control effort for similar attitude errors. In [stingu:09a], the authors describe a linear attitude controller using Euler angle errors for a quadrotor UAV, and the controller works as long as the quadrotor’s attitude is sufficiently far away from the gimbal-lock singularity. In fact, such linear controllers are still prevalent in many commercial jet airplane autopilots which operate for the most part far away from the singular Euler angle attitude. Other similar examples from literature include a design using the truncated Taylor series approximation for the quaternion exponentials in order to achieve a fast single-axis reorientation slew-rate on agile spacecrafts in [gupta:98a], a PID attitude controller in [wie:02a], and an optimal linear controller in [ghiglino:15a].
In this paper, we are interested in analyzing the stability of linear quaternion differential equations in time. In particular we consider the special case of a constant quaternion difference, since the results for this special case are sufficiently novel (the results have not been published previously in the literature to the best of the author’s knowledge) and remarkable (being one of the few problems whose solutions can be analyzed and solved by hand). The chief contributions of this article consist of a novel stability analysis of the constant quaternion difference system, and the presentation of an efficient control design base on that analysis.
@@ II Notation and Terminology
The attitude of a rigid body will be represented by a unit-magnitude 4-component quaternion $\check{q}=[q_{0}\quad q_{1}\quad q_{2}\quad q_{3}]^{\text{T}}$, with a check accent above the symbol representing it. The first component $q_{0}$ of $\check{q}$ is also referred to as the scalar component $\check{q}_{\text{s}}$, while the three components $q_{1}\quad q_{2}\quad q_{3}]^{\text{T}}$ are also referred to as the vector part $\vec{q\mkern-1.0mu}_{\text{v}}$. The most important law of quaternion algebra is their multiplication rule [PhillipsWF:10a]:
$\check{p}\otimes\check{q}$ $=\begin{bmatrix}p_{0}\\ \vec{p\mkern-1.0mu}_{\text{v}}\end{bmatrix}\otimes\begin{bmatrix}q_{0}\\ \vec{q\mkern-1.0mu}_{\text{v}}\end{bmatrix}=\begin{bmatrix}p_{0}q_{0}-\vec{p\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{q\mkern-1.0mu}_{\text{v}}\\ p_{0}\vec{q\mkern-1.0mu}_{\text{v}}+q_{0}\vec{p\mkern-1.0mu}_{\text{v}}+\vec{p\mkern-1.0mu}_{\text{v}}\times\vec{q\mkern-1.0mu}_{\text{v}}\end{bmatrix}$, (1)
where $\vec{p\mkern-1.0mu}_{\text{v}}\times\vec{q\mkern-1.0mu}_{\text{v}}$ (third term in the vector part of the product) denotes the usual vector product of two Euclidean vectors in three-dimensional Euclidean space. The skew-symmetric matrix corresponding to the vector cross product will be denoted as $\vec{p\mkern-1.0mu}_{\text{v}}\times]$ so that $\vec{p\mkern-1.0mu}_{\text{v}}\times\vec{q\mkern-1.0mu}_{\text{v}}$ may also be expressed in matrix notation as $\vec{p\mkern-1.0mu}_{\text{v}}\times]\vec{q\mkern-1.0mu}_{\text{v}}$. It follows from the above multiplication rule that the multiplicative identity is the quaternion with unit scalar part and zero vector part, and that the inverse of a unit-magnitude quaternion is given by negating its vector part [PhillipsWF:10a]:
$\check{q}$ $=\begin{bmatrix}q_{0}\\ \vec{q\mkern-1.0mu}_{\text{v}}\end{bmatrix}\Rightarrow\check{q}^{-1}=\begin{bmatrix}q_{0}\\ -\vec{q\mkern-1.0mu}_{\text{v}}\end{bmatrix}$. (2)
@@
It may be recalled that unit quaternions provide a minimal, singularity-free, and global description of rotations [Shuster:93a], with composition of rotations achieved by multiplication of quaternions. The kinematics of an attitude quaternion may be expressed in terms of the body-frame angular velocity $\vec{\omega\mkern-1.0mu}$ as [PhillipsWF:10a]
$\dot{\check{q}}$ $=\frac{1}{2}\check{q}\otimes\vec{\omega\mkern-1.0mu}.$ (3)
When Euclidean 3-vectors such as $\vec{\omega\mkern-1.0mu}$ are used as quaternion operands as in (3), they are prefixed with a scalar component of zero.
@@
$\vec{n\mkern-1.0mu}$ $=\vec{\omega\mkern-1.0mu}\times J\vec{\omega\mkern-1.0mu}+J\dot{\vec{\omega\mkern-1.0mu}},$ (4)
where $\vec{n\mkern-1.0mu}\in\mathbb{R}^{3}$ is the net external torque, and $J\in\mathbb{R}^{3\times 3}$ is the (second) moment of inertia of the rigid body in the body-fixed coordinate frame. Assuming a solid body, the moment of inertia is positive definite, and the above equation may be inverted to obtain a one-to-one correspondence between the angular acceleration and the applied external torque. Therefore, one may justifiably consider the angular acceleration itself as the user input instead of the torque.
@@
$\dot{\check{p}}$ $=\frac{1}{2}\check{p}\otimes\vec{\nu\mkern-1.0mu},$ (5)
where $\vec{\nu\mkern-1.0mu}$ is the reference angular velocity. We will assume that the angular acceleration $\dot{\vec{\nu\mkern-1.0mu}}$ is also given to us. We will sometimes use the half-angular-velocities, $\vec{v\mkern-1.0mu}$ for $\vec{\nu\mkern-1.0mu}/2$ and $\vec{w\mkern-1.0mu}$ for $\vec{\omega\mkern-1.0mu}/2$, to avoid the factors of 2 and reduce notational clutter. In terms of $\vec{v\mkern-1.0mu}$ and $\vec{w\mkern-1.0mu}$, we have
$\dot{\check{p}}$ $=\check{p}\otimes\vec{v\mkern-1.0mu},\;\dot{\check{q}}=\check{q}\otimes\vec{w\mkern-1.0mu}.$ (6)
III Linear constant difference attitude dynamics
@@ III-A Nominal Dynamics
denote the constant difference between the desired attitude $\check{p}$ and the actual attitude $\check{q}$. Define the unit error quaternion as
@@
$e_{0}$ $=p_{0}q_{0}+\vec{p\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{q\mkern-1.0mu}_{\text{v}}=p_{0}(p_{0}+r_{0})+\vec{p\mkern-1.0mu}_{\text{v}}^{\text{T}}(\vec{p\mkern-1.0mu}_{\text{v}}+\vec{r\mkern-1.0mu}_{\text{v}})$
$=1+r_{0}p_{0}+\vec{r\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{p\mkern-1.0mu}_{\text{v}}.$ (9)
@@
The constant difference constraint (7) may now be invoked to conclude that $e_{0}$, and therefore the error angle, are also constant in time.
The difference quaternion $\check{r}$ is clearly not a unit quaternion in general, but that does not preclude linear control based upon it. We are therefore interested in studying the linear (small perturbation) stability of solutions to the quaternion differential equation
$\dot{\check{q}}=\dot{\check{p}}.$ (12)
Using (6) in (12), we obtain the mathematical statement which encapsulates the constant difference constraint as a relationship between the desired and actual (half) angular velocities, $\vec{v\mkern-1.0mu}$ and $\vec{w\mkern-1.0mu}$:
$\check{p}\otimes\vec{v\mkern-1.0mu}=\check{q}\otimes\vec{w\mkern-1.0mu}\;\Leftrightarrow\;\vec{v\mkern-1.0mu}$ $=\check{e}\otimes\vec{w\mkern-1.0mu},$
@@
$\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{w\mkern-1.0mu}$ $=0,$ (13)
$\vec{v\mkern-1.0mu}$ $=e_{0}\vec{w\mkern-1.0mu}+\vec{e\mkern-1.0mu}_{\text{v}}\times\vec{w\mkern-1.0mu}.$ (14)
@@
It may be noted that we could have as well used the constant difference constraint and the inverse of (8), to infer that $\vec{v\mkern-1.0mu}$ is also perpendicular to $\vec{e\mkern-1.0mu}_{\text{v}}$ like $\vec{w\mkern-1.0mu}$, and to express $\vec{w\mkern-1.0mu}$ in terms of $\vec{v\mkern-1.0mu}$ as follows
$\vec{w\mkern-1.0mu}$ $=\check{e}^{-1}\otimes\vec{v\mkern-1.0mu}$
$\Rightarrow\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{v\mkern-1.0mu}=0,$ (15)
$\hphantom{\Rightarrow}\vec{w\mkern-1.0mu}=e_{0}\vec{v\mkern-1.0mu}-\vec{e\mkern-1.0mu}_{\text{v}}\times\vec{v\mkern-1.0mu}.$ (16)
The former set of equations, (13) and (14), will however turn out to be of greater utility for us in the following stability analysis.
The relationship in (14) may be utilized to eliminate the desired (half) angular velocity $\vec{v\mkern-1.0mu}$ in the error dynamics, which may be derived using (8) and (6) as follows:
$\dot{\check{e}}$ $=-\vec{v\mkern-1.0mu}\otimes\check{p}^{-1}\otimes\check{q}+\check{p}^{-1}\otimes\check{q}\otimes\vec{w\mkern-1.0mu}$
$=\check{e}\otimes\vec{w\mkern-1.0mu}-\vec{v\mkern-1.0mu}\otimes\check{e}.$ (17)
Now, using (14) to substitute for $\vec{v\mkern-1.0mu}$, we may express the kinematics of the error quaternion solely in terms of its own components $e_{0}$ and $\vec{e\mkern-1.0mu}_{\text{v}}$ and the actual (half) angular velocity $\vec{w\mkern-1.0mu}$:
$\dot{\check{e}}$ $=\begin{bmatrix}\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}(\vec{v\mkern-1.0mu}-\vec{w\mkern-1.0mu})\\ e_{0}(\vec{w\mkern-1.0mu}-\vec{v\mkern-1.0mu})+\vec{e\mkern-1.0mu}_{\text{v}}\times(\vec{v\mkern-1.0mu}+\vec{w\mkern-1.0mu})\end{bmatrix}$ (18)
$=\begin{bmatrix}(e_{0}-1)\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{w\mkern-1.0mu}\\ (e_{0}-1)\vec{w\mkern-1.0mu}+\vec{e\mkern-1.0mu}_{\text{v}}\times\vec{w\mkern-1.0mu}+\vec{e\mkern-1.0mu}_{\text{v}}\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{w\mkern-1.0mu}\end{bmatrix}$. (19)
@@
We have so far utilized the relation in (14) to derive the kinematic governing equation (19). The scalar part of (19) is not independent of the vector part since $\check{e}$ is constrained to have unit magnitude, and may be ignored. The vector part of (19) effectively contains three scalar equations in terms of six scalar components, three each in the vectors $\vec{e\mkern-1.0mu}_{\text{v}}$ and $\vec{w\mkern-1.0mu}$. In order to complete the dynamic specification of the system, we need to bring in attitude kinetics and derive an equation for the body (half) angular acceleration $\dot{\vec{w\mkern-1.0mu}}$ which invokes the control torques and moments through (4).
Unlike the dynamics of the error quaternion $\check{e}$, the scalar and vector parts of the dynamics of the difference quaternion $\check{r}$ are independent, since its components are not constrained to yield unit magnitude. Therefore, both (13) and (14) need to be enforced in the nominal motion. We have already exhausted the three scalar degrees of constraint contained in (14) in deriving the three independent scalar equations comprising the vector part of (19). The constraint in (13) has to be incorporated yet, and to this end, we take its derivative to invoke the moment equation, and using (19), we obtain one scalar equation:
$\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\dot{\vec{w\mkern-1.0mu}}=-\vec{w\mkern-1.0mu}^{\text{T}}\dot{\vec{e\mkern-1.0mu}}_{\text{v}}=(1-e_{0})|\vec{w\mkern-1.0mu}|^{2}-(\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{w\mkern-1.0mu})^{2}.$ (20)
We may now design a nominal $\dot{\vec{w\mkern-1.0mu}}$ of minimal magnitude:
$\dot{\vec{w\mkern-1.0mu}}$ $=\frac{(1-e_{0})|\vec{w\mkern-1.0mu}|^{2}\vec{e\mkern-1.0mu}_{\text{v}}}{|\vec{e\mkern-1.0mu}_{\text{v}}|^{2}}=\frac{|\vec{w\mkern-1.0mu}|^{2}\vec{e\mkern-1.0mu}_{\text{v}}}{1+e_{0}}.$ (21)
It may be noted that (21) reflects only the nominal dynamics of the angular velocity. In practice, one usually adds a feedback controller to asymptotically drive the system towards satisfying (13). See Section IV for one possible linear design that yields asymptotic convergence on the angular velocity.
III-B Linearized (Perturbation) Dynamics